OpenStudy (anonymous):

Use any model you choose to solve the story problem. The band director is splitting 1/4 of the band into 4 equal rehearsal groups. What fraction of the entire band will be in each rehearsal group?

OpenStudy (igreen):

1/4 is divided into 4 groups.

OpenStudy (igreen):

Divide 1/4 by 4, can you do that?

OpenStudy (igreen):

Flip the number you're dividing by and multiply. \(\sf\dfrac{1}{4} \div 4 \rightarrow \dfrac{1}{4} \times\dfrac{1}{4}\)